What is the area of of a rectangle that is 10 3/4 inches by 2 3/4 inches?
Answer by Maths68(1474)

Area of rectangle = 10 3/4 * 2 3/4
Area of rectangle = 43/4 * 11/4
Area of rectangle = 473/16
Area of rectangle = 473/16
Area of rectangle = 29 9/16 inches
Area of rectangle = (24+5) 9/16
12 inches = 1 foot
24 inches = 2 feet
Area of rectangle = 2 feet & 5 9/16 inches.